OpenStudy (anonymous):

The graph of -sqrt(x) is shifted 4 units up and 3 units right , what equation represents the new graph..? A. Y=-sqrt(x+4+3) B. Y=-sqrt(x-3+4) C. Y=-sqrt(x+3+4) D. Y=-sqrt(x-4+3)

OpenStudy (anonymous):

None of these is correct. I think you probably meant to type: A. Y = -sqrt(x+4) +3 B. Y = -sqrt(x-3) + 4 etc.

OpenStudy (anonymous):

These are my answer choices

OpenStudy (anonymous):

To shift a curve 3 units to the right, you replace *every* x in the equation by x-3 (NOT x + 3)! When you work with transformations involving x, it goes against your intuition! Shifting a curve up/down is much more intuitive: For example, if you want to shift " y = stuff" up 4, you'd get the new equation "y = stuff + 4". Now you try it...

OpenStudy (anonymous):

(The correct answer is indeed in the image you posted...)

OpenStudy (anonymous):

If you want the full scoop about shifting up/down/left/right, you can look here: http://www.onemathematicalcat.org/Math/Precalculus_obj/shiftUDLR.htm